OpenStudy (anonymous):

Factor the GCF and then the resulting trinomial: 3x3 - 3x2 - 6x x(3x - 6)(x + 1) 3(x - 2)(x + 1) 3x(x - 2)(x + 1) prime

OpenStudy (anonymous):

3(x - 2)(x + 1)

OpenStudy (callisto):

3x3 - 3x2 - 6x = 3x(x^2-x-2) = 3x(x-2)(x+1) can you choose the answer yourself?

OpenStudy (anonymous):

@Callisto you should match with th formulas

OpenStudy (callisto):

@Rohangrr can you explain what formula please?

OpenStudy (anonymous):

@Callisto dude dont take to ur heart i just reminded u the post was mistakenly putted by my bro in my absence

OpenStudy (callisto):

Hmm.. i'm not a guy, so i'm not a dude? and sorry that doesn't sound like 'reminding' me. Anyway, does the asker know how to do it?

OpenStudy (anonymous):

No, then again the comments here are highly entertaining to be honest.

OpenStudy (callisto):

lol, first take out the common factor... you can see that in 3x3 - 3x2 - 6x x is the common factor (as there is an x in each term ) 3 is another common factor (as the coefficients are multiple of 3), understand this part?

OpenStudy (anonymous):

Kind of, the whole factoring part doesn't make sense to me period.

OpenStudy (callisto):

just do it step by step then :)

OpenStudy (callisto):

So, next, take out 3x as the common factor in 3x3 - 3x2 - 6x you'll get 3x(x^2-x-2), understand this part?

OpenStudy (anonymous):

3x^3-3x^2- 6x In all the three terms x is common and 3 is also common so, we can take 3x outside and re-write as 3x(x^2-x-2) (x^2-x-2) can factorise as (x-2)(x+1) so we can factorise as 3x(x-2)(x+1) here, GCF= 3x trinomial can factorise as 3x(x-2)(X+1)